The Tessmer Company has fixed costs of $400,000 and variable costs are 75% of the selling...

  1. The Tessmer Company has fixed costs of $400,000 and variable costs are 75% of the selling price. To realize profits of $100,000 from sales of 500,000 units, the selling price per unit:
  1. Must be $1.33
  2. Must be $4.00
  3. Must be $1.00
  4. Is indeterminable
Answer #1

Let selling price per unit be x

Target sales = (Fixed costs+Target profit)/Contribution margin per unit

500,000 = (400,000+100,000)/(x-0.75x)

500,000 = (500,000)/(0.25x)

0.25x = 500,000/500,000

0.25x = 1

x = 1/0.25

x = 4

Option B is the answer